OpenStudy (anonymous):

Which is equivalent to (3x+5) + 2x = 35? A. 3x+(5+2x)=35 B. 15x=35 C. 8x+2=35 D. 10x=35 person who answers get a fan and medal ^^

OpenStudy (anonymous):

(3x+5) + 2x = 35 3x + 5 + 2x = 35 5x + 5 = 35 5x = 35 - 5 5x = 30 x = 30/5 x = 6

OpenStudy (anonymous):

now check which of the choices yields x=6

OpenStudy (anonymous):

Annas could help xD

OpenStudy (anonymous):

well yep I got it thank yu so much I'll fan everyone who decided to reply xDD thanks ^^

OpenStudy (shadowlegendx):

Annas will now delete all his posts .-.

OpenStudy (anonymous):

*Shadow now replies* Dun dun dunnnnnn

OpenStudy (shadowlegendx):

xD

OpenStudy (anonymous):

lolol

OpenStudy (anonymous):

You are welcome